OpenStudy (anonymous):

HELP! WILL FAN AND MEDAL!!! HELP!!! What is the length of the longest side of a triangle that has the vertices (-3, 2), (3, 2), and (3, 6)? A. B. C. D.

OpenStudy (logan13):

wut u think

OpenStudy (anonymous):

what are the choices

OpenStudy (anonymous):

A. https://media.studyisland.com/cgi-bin/mimetex.cgi?5%20 \%20\text{units} B. https://media.studyisland.com/cgi-bin/mimetex.cgi?2 \sqrt{5}%20\%20\text{units} C. https://media.studyisland.com/cgi-bin/mimetex.cgi? \sqrt{5}%20\%20\text{units} D. https://media.studyisland.com/cgi-bin/mimetex.cgi?2 \sqrt{13}%20\%20\text{units}

OpenStudy (anonymous):

just plot the points and measure the longest side I can't see any of your diagrams

OpenStudy (anonymous):

Copy and paste them? @jamaica25